OpenStudy (anonymous):

@mathstudent55 @amistre64 @ganeshie8 @thomaster For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (amistre64):

can you tell me the definition of closed?

OpenStudy (anonymous):

Not sure what it is asking, but if I am assuming that by closed it means equal to zero, then it would be Addition.

OpenStudy (amistre64):

consider when something is closed, or locked, nothing new can get in, and nothing old can get out .....

OpenStudy (anonymous):

i think it is multipication

OpenStudy (amistre64):

multiplicaiton works

OpenStudy (amistre64):

if we have a closed set: {a,b,c} then for an operation, O, to be closed on the set: xOy is also in the set

OpenStudy (amistre64):

for any x,y in the set

OpenStudy (anonymous):

but it is more than one answer

OpenStudy (perl):

a set is closed under an operation if whenever you operate on two members of the set, the result of the operation is also in the set.

OpenStudy (amistre64):

A. addition 1+1 = 2 B. division 1/0 = ? C. multiplication D. subtraction -1-1 = -2

OpenStudy (amistre64):

it cant be more than one answer ....

OpenStudy (anonymous):

Choose all answers that are correct. it says

OpenStudy (amistre64):

so choose all the answer that are correct

OpenStudy (amistre64):

you dont need to do all the permutation, if one of them fails than its not closed.

OpenStudy (amistre64):

if you cant think of a fail, testing out the permutes is viable

OpenStudy (anonymous):

5000 medals wow @mathstudent55

OpenStudy (anonymous):

so just go with multipication

OpenStudy (amistre64):

yes, its the only one that works

OpenStudy (mathstudent55):

Don't exaggerate; I only have 4604 so far.

OpenStudy (amistre64):

it is ALL the correct ones lol

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

I thought that title was earned at 5000 sorry

OpenStudy (perl):

well 1 + 1 = 2 (unless its modulo 2 addition)

OpenStudy (perl):

only C works , if we have base 10 addition, subtraction, addition, multiplication

OpenStudy (mathstudent55):

@perl You are correct. A number can be paired with itself, so 1 + 1 has to be included. There are 9 operations that need to be checked for each of addition, subtraction, multiplication, and division.

OpenStudy (perl):

the question is a tiny bit ambiguous since it does not state what kind of addition it is (it could be modulo 2 addition) .

OpenStudy (anonymous):

what is modulo 2 addition?

OpenStudy (perl):

thx @mathstudent55

OpenStudy (perl):

modulo addition is like having a clock with only two digits , 0 and 1

OpenStudy (anonymous):

oh I see, so its basically like limitations when graphing?

OpenStudy (perl):



OpenStudy (perl):

you are probably more accustomed to using modulo 12 time

OpenStudy (anonymous):

Ah yes I see now, just wikid it.

OpenStudy (perl):

but it doesn't look like its directly relevant to the question

OpenStudy (anonymous):

true true, what math course uses the modular arithmetic stuff?